# Amplitude Expansion Phase Field Crystal (APFC) Modeling based Efficient Dislocation Simulations using Fourier Pseudospectral Method

### v2

@@ 4.2.1 Density $n$ and Amplitude Parameter $A^{2}$
We employ the same parameters used in Section 4.1 to solve the APFC model (2.13). Based on how we apply the initial values, we study the distribution of the parameter $A^{2}$ defined in (2.16) on the planes $z=L_{z}/4$ and $z=3L_{z}/4$. The distributions on these planes are similar; thus, we present the case for the $z=L_{z}/4$ plane without loss of generality.
Figure 7 visualizes the density $n$ on the $yz$-plane with $x=L_{x}/4$ and the parameter $A^{2}$ on the $z=L_{z}/4$ plane, illustrating the distribution of dislocations as the system stabilizes. Figure 7a and Figure 7b display the density $n$ of the perfect lattice and the defect lattice, which can be reconstructed from (2.2). Similar to the two-dimensional situation, the area near the defect is extracted in Figure 7b, and the burgers vector can be clearly seen. We focus on the regions where $A^{2}$ decreases in Figure 7c. The navy blue area corresponds to the region where $A^{2}<0.8\max(A^{2})$, indicating the presence of defects as $A^{2}$ significantly decreases. The distribution in the figure demonstrates the property of the parameter $A^{2}$: it remains constant within ordered phases but decreases in the presence of defects.
@@
Figure 7: BCC lattice: The illustration of the density $n$ defined in (2.2) for the perfect lattice and defect lattice, and the parameter $A^{2}$ on the $z=L_{z}/4$ plane. Both are colored by their respective values.
4.2.2 Comparison of the Strain Field with Continuum Elasticity Theory
We employ the least squares method to solve the system (4.3). After some straightforward calculations, the least squares solution of (4.3) in the context of the Moore-Penrose inverse is expressed as follows,
$u(\bm{r})=\left\{\begin{array}[]{lr}u_{x}=\frac{\sqrt{2}}{4}(\varphi_{1}+\varphi_{2}+\varphi_{5}-\varphi_{6})\\ u_{y}=\frac{\sqrt{2}}{4}(\varphi_{1}+\varphi_{3}+\varphi_{4}-\varphi_{5})\\ u_{z}=\frac{\sqrt{2}}{4}(\varphi_{2}+\varphi_{3}-\varphi_{4}+\varphi_{6})\end{array}\right..$ (4.14)
Strain Field
According to the shear stress applied in (4.13) and the defect distribution shown in Figure 7, we choose to study the strain field distribution on the $xz$-plane and $yz$-plane. Given the symmetric structure of the BCC lattice, their distributions should be the same. Without loss of generality, we consider the strain field on the $yz$-plane with $x=L_{x}/4$. Similar to the two-dimensional case, the strain field $\varepsilon_{yz}$ can be computed using the following formulation:
@@ 4.2.2 Comparison of the Strain Field with Continuum Elasticity Theory
Figure 8 shows the strain field $\varepsilon_{yz}$ for the three-dimensional BCC lattice. On this plane, we observe parallel dislocations, and the distribution of the field at each defect core aligns with the theoretical expectations discussed in prior work [39].
@@
Comparison with CE
To verify the accuracy of the APFC solution, we examine the strain field $\varepsilon_{yz}$ along the $z=L_{z}/4$ direction on the $yz$-plane and compare the strain field from the APFC model (2.13) with the theoretical results derived from continuum elasticity theory.
In the three-dimensional case, based on the distribution of defects shown in Figure 7, we need to consider two sets of dislocations, with both dislocation cores located at $\bm{x_{c}}=(0,0,0)$ and the Burgers vectors perpendicular to each other, aligned parallel to the $\bm{x}$-axis and $\bm{y}$-axis, respectively. Similar to the two-dimensional case, the strain field $\varepsilon_{yz}$ can be obtained as follows:
$\begin{split}\varepsilon_{yz}&=\frac{1+v}{E}\sigma_{yz}.\end{split}$ (4.15)
Figure 9: BCC lattice: Strain field $\varepsilon_{yz}$ compared with CE along the $z=L_{z}/4$ direction on the $yz$-plane. The red solid lines represent the results from CE theory, while the black circles represent the results from the APFC model (2.13).
As shown in Figure 9, the strain field obtained from the APFC model (2.13) closely matches those derived from continuum elasticity (CE). Notably, there is significant numerical variation around the dislocation cores, which is consistent with the expected strain field behavior in these regions. This outcome reaffirms the effectiveness of the Fourier pseudospectral method in solving the three-dimensional APFC model (2.13), producing results that align well with continuum elasticity theory.
Far Field Decay
@@
Figure 10: BCC lattice: The decay of the strain field $\varepsilon_{yz}$ along $z=L_{z}/4$ direction on the $yz$-plane in terms of the distance $r$ from the dislocation core. The red line represents the results from the CE theory while the green line shows the results by using the APFC model (2.13). Moreover, the blue line represents the error between them.
@@ 5 Conclusion
In this paper, we have demonstrated the Fourier pseudospectral method for solving the APFC model in dislocation simulations. By leveraging the system’s inherent periodicity and applying periodic boundary conditions, the Fourier pseudospectral method offers exceptional accuracy and computational efficiency. Our numerical experiments on two-dimensional triangular lattices and three-dimensional body-centered cubic lattices successfully optimized edge dislocation geometries, yielding strain fields consistent with continuum elasticity theory predictions. These results highlight the potential of this method in advancing the simulation of complex crystalline defects. Our findings suggest broad applicability of this numerical approach to other crystalline defects, several open problems warrant further investigation:
@@
  • Closing the gap between the APFC and atomistic models: Aligning the elastic parameters of the APFC and atomistic models is key to establishing a connection between them. Future research should explore coupling methods similar to classical atomistic-to-continuum techniques to effectively bridge these models [50].
@@
Acknowledgements
XW, YW and LZ are partially supported by NSFC grant 12271360. KJ is partially supported by the National Key R&D Program of China 2023YFA1008802, the National Natural Science Foundation of China 12171412, the Science and Technology Innovation Program of Hunan Province 2024RC1052 and the Innovative Research Group Project of Natural Science Foundation of Hunan Province of China 2024JJ1008. YW is also partially supported by the Development Postdoctoral Scholarship for Outstanding Doctoral Graduates from Shanghai Jiao Tong University. LZ is also partially supported by the Shanghai Municipal Science and Technology Project 22JC1401600 and the Fundamental Research Funds for the Central Universities.
Appendix A Introduction to Bravais Lattice and Reciprocal Lattice
In this section, we provide an introduction to Bravais lattices and reciprocal lattices, as well as their mathematical formulations. For any $d$-dimensional periodic structure $F(\bm{r})$, where $\bm{r}\in\mathbb{R}^{d}$, the repeated structural unit is called a unit cell. A fundamental unit cell $\Omega$ is described by $d$ vectors $(\bm{e_{1}},\bm{e_{2}},\ldots,\bm{e_{d}})$:
$\Omega=\left\{x\in\Omega\,\bigg{|}\,x=\sum_{j=1}^{n}\alpha_{j}\bm{e_{j}},\,\alpha_{j}\in[0,1]\right\}.$ (A.1)
@@ Appendix A Introduction to Bravais Lattice and Reciprocal Lattice
$\bm{R_{n}}=n_{1}\bm{e_{1}}+n_{2}\bm{e_{2}}+\ldots+n_{d}\bm{e_{d}},\quad n_{i}\in\mathbb{Z},$
@@
The reciprocal space lattice is given by:
$\bm{G_{m}}=m_{1}\bm{e_{1}}^{*}+m_{2}\bm{e_{2}}^{*}+\ldots+m_{d}\bm{e_{d}}^{*},\quad m_{i}\in\mathbb{Z},$
@@
One of the most important properties of the reciprocal lattice vector is that the plane wave $\{e^{\imath\bm{G_{m}}\cdot\bm{r}}\}$ can form a set of basis functions for any periodic function $f\in L^{2}(\mathbb{T}_{N}^{d})$ [51]. This can be expressed as:
$f(\bm{r})=\sum_{\bm{m}}\widehat{f}_{\bm{m}}e^{\imath{\bm{G_{m}}}\cdot{\bm{r}}}.$ (A.2)
For any $\bm{R_{n}}$ in the Bravais lattice, the structure is invariant under lattice translation:
$f(\bm{r}+\bm{R_{n}})=f(\bm{r}).$ (A.3)
Expanding the functions on both sides of equation (A.3) as in equation (A.2), we have:
$\sum_{\bm{m}}\widehat{f}_{\bm{m}}e^{\imath{\bm{G_{m}}}\cdot{\bm{r}}}=\sum_{\bm{m}}\widehat{f}_{\bm{m}}e^{\imath{\bm{G_{m}}}\cdot(\bm{r}+\bm{R_{n}})}=\sum_{\bm{m}}\widehat{f}_{\bm{m}}e^{\imath{\bm{G_{m}}}\cdot{\bm{r}}}e^{\imath{\bm{G_{m}}}\cdot{\bm{R_{n}}}}.$
Note that $e^{\imath{\bm{G_{m}}}\cdot{\bm{R_{n}}}}=1$ holds only if the following equation is satisfied:
$\bm{G_{m}}\cdot\bm{R_{n}}=2\pi N,\quad N\in\mathbb{Z}.$ (A.4)
Hence, equation (A.2) can be rewritten as:
$f(\bm{r})$ $=\sum_{\bm{m}}\widehat{f}_{\bm{m}}e^{\imath{\bm{G_{m}}}\cdot{\bm{r}}}$ (A.5)
$=\sum_{\bm{k}\in K_{N}^{n}}\widehat{f}_{\bm{k}}e^{\imath\left(\frac{2\pi}{\bm{R_{n}}}\cdot\bm{k}\right)\cdot{\bm{r}}}$
$=\sum_{\bm{k}\in K_{N}^{n}}\widehat{f}_{\bm{k}}e^{\imath{(B\bm{k})\cdot{\bm{r}}}},$
where $B=\frac{2\pi}{\bm{R_{n}}}$ is the transformation matrix from physical space to reciprocal space.
@@ Appendix B Auxiliary Results
B.1 The derivation of (2.14)
We first derive (2.14). We start by calculating $A^{2}$ and $A^{4}$:
$A^{2}$ $=2\sum_{j=1}^{N}|\eta_{j}|^{2}=2\sum_{j=1}^{N}\eta_{j}\eta_{j}^{*},$
$A^{4}$ $=4\left(\sum_{j=1}^{N}|\eta_{j}|^{2}\right)^{2}=4\left(\sum_{j=1}^{N}\eta_{j}\eta_{j}^{*}\right)^{2}.$
Substituting these into (2.15), we obtain
$F_{\eta_{j}}$ $=\frac{1}{|\Omega|}\int_{\Omega}\Big{[}\frac{B_{0}}{2}A^{2}+\frac{3v}{4}A^{4}+\sum_{j=1}^{N}\left(B_{x}|\varsigma_{j}\eta_{j}|^{2}-\frac{3v}{2}|\eta_{j}|^{4}\right)+f^{s}({\eta_{j}},{\eta_{j}^{*}})\Big{]}\,{\rm d}\bm{r}$ (B.1)
$=\frac{1}{|\Omega|}\int_{\Omega}\Big{[}B_{0}\sum_{j=1}^{N}|\eta_{j}|^{2}+3v\left(\sum_{j=1}^{N}\eta_{j}\eta_{j}^{*}\right)^{2}+\sum_{j=1}^{N}\left(B_{x}|\varsigma_{j}\eta_{j}|^{2}-\frac{3v}{2}|\eta_{j}|^{4}\right)+f^{s}({\eta_{j}},{\eta_{j}^{*}})\Big{]}\,{\rm d}r.$
The $i$-th partial functional derivative $\frac{\delta F}{\delta\eta_{j}^{*}}$ of $F$ with respect to $\eta_{j}$ can be calculated as follows. For any test function $\phi_{j}\in U$, where $U$ is a Banach space of functions $\eta_{j}$ on the computational domain $\Omega\in\mathbb{R}^{d}$, we have
$\left(\frac{\delta F}{\delta\eta_{j}^{*}},\phi_{j}\right)$ $=\frac{d}{dt}F(\eta_{1},\ldots,\eta_{N},\eta_{1}^{*},\ldots,\eta_{N}^{*})\bigg{|}_{t=0}$
$=\lim_{t\to 0}\frac{F(\eta_{1},\ldots,\eta_{N},\eta_{1}^{*},\ldots,\eta_{j}^{*}+t\phi_{j},\ldots,\eta_{N}^{*})-F(\eta_{1},\ldots,\eta_{N},\eta_{1}^{*},\ldots,\eta_{N}^{*})}{t}$
$=\lim_{t\to 0}\frac{1}{t|\Omega|}\int_{\Omega}\left[B_{0}t\eta_{j}\phi_{j}+3v\left((A^{2}-\eta_{j}\eta_{j}^{*})+\eta_{j}t\phi_{j}\right)(t\eta_{j}\phi_{j})+B_{x}t\varsigma_{j}^{2}\eta_{j}\phi_{j}-\frac{3v}{2}t\eta_{j}^{2}\phi_{j}^{2}\right]\,{\rm d}\bm{r}+\left(\frac{\delta f^{s}}{\delta\eta_{j}^{*}},\phi_{j}\right)$
$=\frac{1}{|\Omega|}\int_{\Omega}\left[B_{0}\eta_{j}\phi_{j}+3v\left(A^{2}-|\eta_{j}|^{2}\right)\eta_{j}\phi_{j}+B_{x}\varsigma_{j}^{2}\eta_{j}\phi_{j}\right]\,{\rm d}\bm{r}+\left(\frac{\delta f^{s}}{\delta\eta_{j}^{*}},\phi_{j}\right)$
$=\left(\left[B_{0}+3v\left(A^{2}-|\eta_{j}|^{2}\right)+B_{x}\varsigma_{j}^{2}\right]\eta_{j},\phi_{j}\right)+\left(\frac{\delta f^{s}}{\delta\eta_{j}^{*}},\phi_{j}\right).$

### v3

@@ 4.2.1 Density $n$ and amplitude parameter $A^{2}$
We employ the same parameters used in Section 4.1 to solve the APFC model (13). Based on how we apply the initial values, we study the distribution of the parameter $A^{2}$ defined in (16) on the planes $z=L_{z}/4$ and $z=3L_{z}/4$. The distributions on these planes are similar; thus, we present the case for the $z=L_{z}/4$ plane without loss of generality.
Figure 9 visualizes the density $n$ on the $yz$-plane with $x=L_{x}/4$ and the parameter $A^{2}$ on the $z=L_{z}/4$ plane, illustrating the distribution of dislocations as the system stabilizes. Figure 9a and Figure 9b display the density $n$ of the perfect lattice and the defect lattice, which can be reconstructed from (2). Similar to the two-dimensional situation, the area near the defect is extracted in Figure 9b, and the Burgers vector can be clearly seen. We focus on the regions where $A^{2}$ decreases in Figure 9c. The navy blue area corresponds to the region where $A^{2}<0.8\max(A^{2})$, indicating the presence of defects as $A^{2}$ significantly decreases. The distribution in the figure demonstrates the property of the parameter $A^{2}$: it remains constant within ordered phases but decreases in the presence of defects.
@@
Figure 9: BCC lattice: The illustration of the density $n$ defined in (2) for the perfect lattice and defect lattice, and the parameter $A^{2}$ on the $z=L_{z}/4$ plane. Both are colored by their respective values.
4.2.2 Strain field and comparison with continuum elasticity
We employ the least squares method to solve the system (33). After some straightforward calculations, the least squares solution of (33) in the context of the Moore-Penrose inverse is expressed as follows:
$\bm{u}(\bm{r})=\left\{\begin{array}[]{lr}u_{x}=\frac{\sqrt{2}}{4}(\varphi_{1}+\varphi_{2}+\varphi_{5}-\varphi_{6})\\ u_{y}=\frac{\sqrt{2}}{4}(\varphi_{1}+\varphi_{3}+\varphi_{4}-\varphi_{5})\\ u_{z}=\frac{\sqrt{2}}{4}(\varphi_{2}+\varphi_{3}-\varphi_{4}+\varphi_{6})\end{array}\right.$. (44)
Strain field.
According to the shear stress applied in (43) and the defect distribution shown in Figure 9, we choose to study the strain field distribution on the $xz$-plane and $yz$-plane. Given the symmetric structure of the BCC lattice, their distributions should be the same. Without loss of generality, we consider the strain field on the $yz$-plane with $x=L_{x}/4$. Similar to the two-dimensional case, the strain field $\varepsilon_{yz}$ can be computed using the following formulation:
@@ 4.2.2 Strain field and comparison with continuum elasticity
Figure 10 shows the strain field $\varepsilon_{yz}$ for the three-dimensional BCC lattice. On this plane, we observe parallel dislocations, and the distribution of the field at each defect core aligns with the theoretical expectations discussed in prior work anderson2017theory .
@@
Comparison with continuum elasticity theory.
To verify the accuracy of the APFC solution, we examine the strain field $\varepsilon_{yz}$ along the $z=L_{z}/4$ direction on the $yz$-plane and compare the strain field from the APFC model (13) with the theoretical results derived from continuum elasticity theory.
In the three-dimensional case, based on the distribution of defects shown in Figure 9, we need to consider two sets of dislocations, with both dislocation cores located at $\bm{x}_{c}=(0,0,0)$ and the Burgers vector perpendicular to each other, aligned parallel to the $\bm{x}$-axis and $\bm{y}$-axis, respectively. Similar to the two-dimensional case, the strain field $\varepsilon_{yz}$ can be obtained as follows:
$\begin{split}\varepsilon_{yz}&=\frac{1+v}{E}\sigma_{yz}.\end{split}$ (45)
Figure 11: BCC lattice: Strain field $\varepsilon_{yz}$ compared with CE along the $z=L_{z}/4$ direction on the yzyz-plane. The red solid lines represent the results from CE theory, while the black circles represent the results from the APFC model (13).
As shown in Figure 11, the strain field obtained from the APFC model (13) closely matches those derived from continuum elasticity (CE). Notably, there is significant numerical variation around the dislocation cores, which is consistent with the expected strain field behavior in these regions. This outcome reaffirms the effectiveness of the Fourier pseudospectral method in solving the three-dimensional APFC model (13), producing results that align well with continuum elasticity theory.
Far-field decay.
@@
Figure 12: BCC lattice: The decay of the strain field $\varepsilon_{yz}$ along $z=L_{z}/4$ direction on the $yz$-plane in terms of the distance $r$ from the dislocation core. The red line represents the results from the CE theory while the green line shows the results by using the APFC model (13). Moreover, the blue line represents the error between them.
@@ 5 Conclusion
In this paper, we have demonstrated the Fourier pseudospectral method for solving the APFC model in dislocation simulations. By leveraging the system’s inherent periodicity and applying periodic boundary conditions, the Fourier pseudospectral method offers exceptional accuracy and computational efficiency. Our numerical experiments on two-dimensional triangular lattices and three-dimensional body-centered cubic lattices successfully optimized edge dislocation geometries, yielding strain fields consistent with the predictions of continuum elasticity theory. These results highlight the potential of this method in advancing the simulation of complex crystalline defects. Our findings suggest broad applicability of this numerical approach to other crystalline defects. Several open problems warrant further investigation:
@@
  • Closing the gap between the APFC and atomistic models: Aligning the elastic parameters of the APFC and atomistic models is key to establishing a connection between them. Future research should explore coupling methods similar to classical atomistic-to-continuum techniques to effectively bridge these models 2019Closing .
@@
Appendix A Introduction to Bravais Lattices and Reciprocal Lattices
In this section, we provide an introduction to Bravais lattices and reciprocal lattices, as well as their mathematical formulations. For any $d$-dimensional periodic structure $F(\bm{r})$, where $\bm{r}\in\mathbb{R}^{d}$, the repeated structural unit is called a unit cell.
@@ Appendix A Introduction to Bravais Lattices and Reciprocal Lattices
$\bm{R}_{\bm{n}}=n_{1}\bm{e}_{1}+n_{2}\bm{e}_{2}+\ldots+n_{d}\bm{e}_{d},\quad n_{i}\in\mathbb{Z},$
@@
The reciprocal lattice vector is given by:
$\bm{G}_{\bm{m}}=m_{1}\bm{e}_{1}^{*}+m_{2}\bm{e}_{2}^{*}+\ldots+m_{d}\bm{e}_{d}^{*},\quad m_{i}\in\mathbb{Z},$
@@
One of the most important properties of the reciprocal lattice vector is that the plane wave $\{e^{\imath\bm{G}_{\bm{m}}\cdot\bm{r}}\}$ can form a set of basis functions for any periodic function $f\in L^{2}(\mathbb{T}_{N}^{d})$. This can be expressed as
$f(\bm{r})=\sum_{\bm{m}}\widehat{f}_{\bm{m}}e^{\imath{\bm{G}_{\bm{m}}}\cdot{\bm{r}}},$ (46)
where $\widehat{f}_{\bm{m}}$ are Fourier coefficients. For any $\bm{R}_{\bm{n}}$ in the Bravais lattice, the structure is invariant under lattice translation:
$f(\bm{r}+\bm{R}_{\bm{n}})=f(\bm{r}).$ (47)
Expanding the functions on both sides of (47) as in (46), we have
$\sum_{\bm{m}}\widehat{f}_{\bm{m}}e^{\imath{\bm{G}_{\bm{m}}}\cdot{\bm{r}}}=\sum_{\bm{m}}\widehat{f}_{\bm{m}}e^{\imath{\bm{G}_{\bm{m}}}\cdot(\bm{r}+\bm{R}_{\bm{n}})}=\sum_{\bm{m}}\widehat{f}_{\bm{m}}e^{\imath{\bm{G}_{\bm{m}}}\cdot{\bm{r}}}e^{\imath{\bm{G}_{\bm{m}}}\cdot{\bm{R}_{\bm{n}}}}.$
Note that $e^{\imath{\bm{G}_{\bm{m}}}\cdot{\bm{R}_{\bm{n}}}}=1$ holds only if the following equation is satisfied:
$\bm{G}_{\bm{m}}\cdot\bm{R}_{\bm{n}}=2\pi N,\quad N\in\mathbb{Z}.$ (48)
Hence, equation (46) can be rewritten as:
$f(\bm{r})$ $=\sum_{\bm{m}}\widehat{f}_{\bm{m}}e^{\imath{\bm{G}_{\bm{m}}}\cdot{\bm{r}}}$ (49)
$=\sum_{\bm{k}\in K_{N}^{d}}\widehat{f}_{\bm{k}}e^{\imath\left(\frac{2\pi}{\bm{R}_{\bm{n}}}\cdot\bm{k}\right)\cdot{\bm{r}}}$
$=\sum_{\bm{k}\in K_{N}^{d}}\widehat{f}_{\bm{k}}e^{\imath{(B\bm{k})\cdot{\bm{r}}}},$
where $B=2\pi/\bm{R}_{\bm{n}}$ denotes the transformation matrix from physical space to reciprocal space.
@@ Appendix B Auxiliary Results
B.1 The derivation of (14)
We first derive (14). We begin by computing $A^{2}$ and $A^{4}$:
$A^{2}$ $=2\sum_{j=1}^{M}\lvert\eta_{j}\rvert^{2}=2\sum_{j=1}^{M}\eta_{j}\eta_{j}^{*},$
$A^{4}$ $=4\left(\sum_{j=1}^{M}\lvert\eta_{j}\rvert^{2}\right)^{2}=4\left(\sum_{j=1}^{M}\eta_{j}\eta_{j}^{*}\right)^{2}.$
Substituting these into (15), we obtain
$F$ $=\frac{1}{\lvert\Omega\rvert}\int_{\Omega}\Big[\frac{B_{0}}{2}A^{2}+\frac{3v}{4}A^{4}+\sum_{j=1}^{M}\left(B_{x}\lvert\varsigma_{j}\eta_{j}\rvert^{2}-\frac{3v}{2}\lvert\eta_{j}\rvert^{4}\right)+f^{s}({\eta_{j}},{\eta_{j}^{*}})\Big]\,{\rm d}\bm{r}$ (50)
$=\frac{1}{\lvert\Omega\rvert}\int_{\Omega}\Big[B_{0}\sum_{j=1}^{M}\lvert\eta_{j}\rvert^{2}+3v\left(\sum_{j=1}^{M}\eta_{j}\eta_{j}^{*}\right)^{2}+\sum_{j=1}^{M}\left(B_{x}\lvert\varsigma_{j}\eta_{j}\rvert^{2}-\frac{3v}{2}\lvert\eta_{j}\rvert^{4}\right)\Big]\,{\rm d}\bm{r}$
$\quad+\frac{1}{\lvert\Omega\rvert}\int_{\Omega}f^{s}({\eta_{j}},{\eta_{j}^{*}})\,{\rm d}\bm{r}.$
The partial functional derivative $\delta F/\delta\eta_{j}^{*}$ of $F$ with respect to $\eta_{j}$ can be calculated as follows. For any test function $\phi_{j}\in U$, where $U$ is a Banach space of functions $\eta_{j}$ on the computational domain $\Omega\in\mathbb{R}^{d}$, we have
$\left(\frac{\delta F}{\delta\eta_{j}^{*}},\phi_{j}\right)$ $=\frac{d}{dt}F(\eta_{1},\ldots,\eta_{M},\eta_{1}^{*},\ldots,\eta_{j}^{*}+t\phi_{j},\ldots,\eta_{M}^{*})\bigg\lvert_{t=0}$
$=\lim_{t\to 0}\frac{F(\eta_{1},\ldots,\eta_{M},\eta_{1}^{*},\ldots,\eta_{j}^{*}+t\phi_{j},\ldots,\eta_{M}^{*})}{t}$
$\quad-\lim_{t\to 0}\frac{F(\eta_{1},\ldots,\eta_{M},\eta_{1}^{*},\ldots,\eta_{M}^{*})}{t}$
$=\lim_{t\to 0}\frac{1}{t\lvert\Omega\rvert}\int_{\Omega}\left[B_{0}t\eta_{j}\phi_{j}+3v\left((A^{2}-\eta_{j}\eta_{j}^{*})+\eta_{j}t\phi_{j}\right)(t\eta_{j}\phi_{j})\right]\,{\rm d}\bm{r}$
$\quad+\lim_{t\to 0}\frac{1}{t\lvert\Omega\rvert}\int_{\Omega}B_{x}t\varsigma_{j}^{2}\eta_{j}\phi_{j}-\frac{3v}{2}t\eta_{j}^{2}\phi_{j}^{2}\,{\rm d}\bm{r}+\left(\frac{\delta f^{s}}{\delta\eta_{j}^{*}},\phi_{j}\right)$
$=\frac{1}{\lvert\Omega\rvert}\int_{\Omega}\left[B_{0}\eta_{j}\phi_{j}+3v\left(A^{2}-\lvert\eta_{j}\rvert^{2}\right)\eta_{j}\phi_{j}+B_{x}\varsigma_{j}^{2}\eta_{j}\phi_{j}\right]\,{\rm d}\bm{r}+\left(\frac{\delta f^{s}}{\delta\eta_{j}^{*}},\phi_{j}\right)$
$=\left(\left[B_{0}+3v\left(A^{2}-\lvert\eta_{j}\rvert^{2}\right)+B_{x}\varsigma_{j}^{2}\right]\eta_{j},\phi_{j}\right)+\left(\frac{\delta f^{s}}{\delta\eta_{j}^{*}},\phi_{j}\right).$